OpenStudy (anonymous):

if f''(x) exist on [a,b] and f(a)=f(b) . prove that f((a+b)/2)=(f(a)+f(b))/2 +((b-a)^2)*f''(c). for c belongs to (a,b).

OpenStudy (anonymous):

without writing it i will bet that it is mvt (or rolle's theorem) twice yes?

OpenStudy (jamesj):

Are you sure you have written this expression down correctly? For example the term (f(a)+f(b))/2 = f(a) = f(b) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

plz help...

OpenStudy (zarkon):

Let \[f(x)=x^2\]\[a=-1,b=1\] then \[f(a)=f(-1)=1=f(1)=f(b)\] \[f\left(\frac{a+b}{2}\right)=f\left(\frac{-1+1}{2}\right)=f(0)=0^2=0\] \[\frac{f(a)+f(b)}{2}=\frac{1+1}{2}=1\] \[f''(c)=2\text{ for all c}\] then using your formula we get \[0=1+(1-(-1))^2\cdot 2\Rightarrow0=9\]

OpenStudy (anonymous):

but its a particular case....i need a general ans..

OpenStudy (zarkon):

it is a counter example

OpenStudy (jamesj):

Pearls my friend, pearls.

OpenStudy (jamesj):

@jyoti: What Zarkon is saying is that the relation you have written down cannot be true in general because he has just given us an example of where it is false. So either you have written down the question incorrectly here in open study; or, the question as given to you is simply incorrect.

OpenStudy (anonymous):

may be the question is incorrect....bcz i have written exactly what was given.....